# presep= ]e1,e2,e3

### v1

@@ Abstract
We study the consistency and weak convergence of the conditional tail function and conditional Hill estimators under broad dependence assumptions for a heavy-tailed response sequence and a covariate sequence. Consistency is established under $\alpha$-mixing, while asymptotic normality follows from $\beta$-mixing and second-order conditions. A key aspect of our approach is its versatile functional formulation in terms of the conditional tail process. Simulations demonstrate its performance across dependence scenarios. We apply our method to extreme event modeling in the oil industry, revealing distinct tail behaviors under varying conditioning values.
62G32,
62E20, 62M10,
Extreme values,
Time series,
Conditional regular variation,
@@ 1 Introduction
Functional limit theorems for the tail empirical process under dependence conditions have been studied in the literature. Early results were established by Rootzén (1995) for $\alpha$-mixing sequences and later extended by Rootzén (2009) to more general dependence structures. In Drees (2000) the asymptotic normality of the POT estimator was shown under $\beta$-mixing conditions. Among the most widely used estimators for heavy-tailed data is the Hill estimator, originally introduced by Hill (1975). Its consistency in dependent settings was first investigated by Resnick and Stărică (1995), who analysed its behaviour under various dependence structures. Further work by Resnick and Stărică (1997) established the consistency of the tail empirical measure for autoregressive processes with heavy-tailed innovations. The asymptotic normality of the Hill estimator for i.i.d. data has been well documented, with significant contributions from Hall (1982), Davis and Resnick (1984), Haeusler and Teugels (1985), and Csörgő and Mason (1985). Extending these results to stationary sequences, Hsing (1991) proved a central limit theorem for the Hill estimator under $\beta$-mixing conditions, and a modern treatment can be found in Kulik and Soulier (2020).
@@
Other work is related to ours. In the presence of covariates, the asymptotic properties of extreme value estimators have been investigated using regression-based approaches. In Daouia et al. (2010) tail index estimation is studied under $\alpha$-mixing conditions using a smoothed Pickands estimator, a technique later generalized by Daouia, Gardes and Girard (2013); Daouia, Stupfler and Usseglio-Carlevé (2023) to regression models within the general max-domain of attraction. Furthermore, tail index regression in a random covariate setting has been analysed using a maximum likelihood framework, as explored by Wang and Tsai (2009). Finally, asymptotic normality of the conditional kernel Hill estimator has been explored for a special second-order regularly varying class in the preprint Goegebeur and Guillou (2024).
@@ 2.1 Notation and setting
Let $x\in\mathbb{R}$ with $g(x)>0$ be given. Our key assumption is that the conditional distribution of $Y_{0}$, given $X_{0}=x$, is regularly varying with tail index $1/\gamma(x)=\alpha_{x}$, and that this holds in a neighbourhood around $x$, that is
@@
Let $K$ be a bounded density, denoted a Kernel, supported on $[-1,1]$ and denote by $\left\{h_{n},n\in\mathbb{N}\right\}$ the corresponding bandwidth (scale) parameter. Let $\left\{u^{x}_{n},n\in\mathbb{N}\right\}$ be a scaling sequence diverging to infinity. Henceforth, we omit the dependence on $x$ and simply write $\left\{u_{n}\right\}$. The conditional survival function $\overline{F}^{x}$ is estimated with the Nadaraya–Watson pathwise estimator as
$\overline{F}^{x}_{n}(y)=\sum_{j=1}^{n}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)}{\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)}1_{\left\{Y_{j}>y\right\}},\quad y>0.$ (2)
@@
$\widetilde{T}^{x}_{n}(s)=\frac{\overline{F}^{x}_{n}(su_{n})}{\overline{F}^{x}(u_{n})}=\frac{\frac{1}{nh_{n}\overline{F}^{x}(u_{n})}\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}\right\}}}{\frac{1}{nh_{n}}\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)}=:\frac{\widetilde{V}^{x}_{n}(s)}{g_{n}(x)},$ (3)
for $s>0$. The target of $\widetilde{T}^{x}_{n}$ is the conditional tail function $s\mapsto T^{x}(s)=s^{-\alpha_{x}}$. Indeed, this convergence, both in probability and weakly, will be shown under suitable conditions in this paper. However, the deterministic thresholds $u_{n}$ are not observable. Below we provide asymptotic results also for when these thresholds are estimated from data.
@@ 2.2 Conditional tail function and conditional Hill estimators
$\gamma(x)=\int_{1}^{\infty}\frac{T^{x}(s)}{s}\mathrm{d}s.$
@@
Let $u_{n}=F^{\leftarrow,x}\left(1-k_{n}/n\right)$ be an unobserved sequence of deterministic thresholds, diverging to infinity, where $F^{\leftarrow,x}$ denotes the generalised inverse of $F^{x}$. We may clearly estimate this sequence by its empirical counterpart $\left\{q_{n},n\in\mathbb{N}\right\}$, suppressing dependence on $x$, given by
$q_{n}=F^{\leftarrow,x}_{n}\left(1-k_{n}/n\right),\quad n\in\mathbb{N}.$
@@
$\widehat{T}^{x}_{n}(s)=\frac{\overline{F}^{x}_{n}(sq_{n})}{\overline{F}^{x}(u_{n})}=\widetilde{T}^{x}_{n}\left(s\frac{q_{n}}{u_{n}}\right),$
which no longer depends on the sequence $u_{n}$, which can be seen from the definition of $\widetilde{T}^{x}_{n}$ and from the fact that $\overline{F}^{x}(u_{n})=k_{n}/n$.
@@
$\widehat{\gamma}_{n}\left(x\right)=\int_{1}^{\infty}\frac{\widehat{T}_{n}^{x}(s)}{s}\mathrm{d}s=-\int_{1}^{\infty}\log\left(s\right)\widehat{T}_{n}^{x}(\mathrm{d}s)=\frac{n}{k_{n}}\frac{\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{j}}{q_{n}}\right)}{\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)}.$
@@
In case of $K$ being the uniform density on $[-1,1]$, then $\widehat{\gamma}_{n}\left(x\right)$ reduces to the Hill estimator applied to the subsample of the data having $X_{i}\in[x-h_{n},x+h_{n}]$, except that the $k_{n}$ sequence is mapped to the number of top order statistics in the subsample, $\tilde{k}_{n}$, through the following relation:
@@
[…]
Before delving into the mathematical treatment of our estimators, we provide an example of a conditional Hill plot. In Figure 1 we have taken $\left\{Y_{j}\right\}$ to be the (absolute value of) the negative log-returns of the West Texas Intermediate (WTI) grade crude oil spot prices from $2007$-$2024$. As a covariate $\left\{X_{j}\right\}$, we take the Crude Oil Volatility Index (CBOE OVX), transformed to have uniform margins. For large $x$, we clearly observe a different, heavier tail regime. In contrast, for medium and small $x$, the tail regime is similar and lighter.
@@ 3 Main results
In this section we provide consistency and weak convergence results for the estimators of the conditional tail function and conditional tail index under suitable conditions. Most conditions are versions of their most general unconditional counterparts. Notably, only $\alpha$-mixing is required for consistency, while $\beta$-mixing is required for normality. Furthermore, second order conditions are formulated in a very general and abstract form. It is later seen that broad classes of regularly-varying tails satisfy them.
@@ 3.1 Consistency
    The map $z\mapsto\alpha_{z}$ is $c_{\alpha}$-Lipschitz continuous,
@@
    $\lim_{n\to\infty}nh_{n}\overline{F}^{x}(u_{n})=\infty$,
@@
    For any $x\in\mathbb{R}$, it holds $g_{n}(x)\overset{\mathrm{Pr}}{\to}g(x)$, see Lemma 12, 13 or 14,
@@
    $\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}\right\}}1_{\left\{Y_{j}>tu_{n}\right\}}\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n})}\to M^{x}(s,t),$
@@
$\widetilde{T}^{x}_{n}(s)\overset{\mathrm{Pr}}{\to}T^{x}(s)=s^{-\alpha_{x}},\ s>0,$ (4)
@@
If there exists $c_{x}\in\mathbb{R}$ such that
$\frac{\overline{F}^{x}(u_{n})+h_{n}^{c_{x}\eta-2}}{nh_{n}^{c_{x}}(\overline{F}^{x}(u_{n}))^{2}}\to 0,$ (5)
@@
As a first application we get that consistency of the conditional tail empirical function implies that the intermediate empirical quantile $q_{n}$ is equivalent to the deterministic $u_{n}$. This is of practical importance, since the former is estimable from data. Consistency of the conditional Hill estimator then directly follows.
@@
If (4) holds, then $q_{n}/u_{n}\overset{\mathrm{Pr}}{\to}1.$
@@
Assume conditions 1-3, 5 and (4). Then $\widehat{\gamma}_{n}\left(x\right)\overset{\mathrm{Pr}}{\to}\gamma(x)$.
@@ 3.2 Central limit theorems
Notice that the speed of convergence is determined by the effective sample size $nh_{n}\overline{F}^{x}(u_{n})=h_{n}k_{n}$, which is slower than in the unconditional case, which is simply $k_{n}$. This price to pay, however, results in a variance which agrees with the i.i.d. conditional case, in contrast to the more involved expression for unconditional time series provided as an infinite sum, e.g. in Kulik and Soulier (2020, Eq. (9.5.4)).
@@
$B_{n}^{x}(s_{0})=\sup_{t\geq s_{0}}\left|\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-\alpha_{x}}g(x)\right|.$
@@
    For $s,t>0$,
    $\lim_{m\to\infty}\limsup_{n\to\infty}\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n})}=0,$
@@
    For $A>0$,
    $\lim_{m\to\infty}\limsup_{n\to\infty}\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}A}\right)\right]}{h_{n}\overline{F}^{x}(u_{n})}=0,$
@@
    $\lim_{n\to\infty}\sqrt{k_{n}h_{n}}\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-\alpha_{x}}g(x)}{t}\mathrm{d}t=0.$
@@
$\Gamma(s,t)=\frac{\left(s\vee t\right)^{-\alpha_{x}}}{g(x)}\int K^{2}(u)\mathrm{d}u,$ (6)
@@
[…]
@@
[…]
The sequence $\{u_{n}\}$ is unobserved and we may estimate it with $\{q_{n}\}$. Let
@@
[…]
@@
[…]
@@
provided conditions involving the logarithmic function are replaced with suitably modified versions involving $\phi$. In case the functional $z\mapsto\int\phi\mathrm{d}y$, or any other functional is defined on $\mathbb{D}\left([A,B]\right)$ and is sup-norm continuous or Hadamard differentiable, consistency and a central limit theorem for the associated estimator follow directly from Theorem 4, together with the continuous mapping theorem or functional delta-method, respectively.

### v2

# Conditional Extreme Value Estimation for Dependent Time Series

@@ Abstract
We study the consistency and weak convergence of the conditional tail function and conditional Hill estimators under broad dependence assumptions for a heavy-tailed response sequence and a covariate sequence. Consistency is established under $\alpha$-mixing, while asymptotic normality follows from $\beta$-mixing and second-order conditions. A key aspect of our approach is its versatile functional formulation in terms of the conditional tail process. Simulations demonstrate its performance across dependence scenarios. We apply our method to extreme event modelling in the oil industry, revealing distinct tail behaviours under varying conditioning values.
@@ 1 Introduction
Functional limit theorems for the tail empirical process under dependence conditions have been studied in the literature. Early results were established by Rootzen_1995 for $\alpha$-mixing sequences and later extended by Rootzen_2009 to more general dependence structures. In Drees_2000 the asymptotic normality of the POT estimator was shown under $\beta$-mixing conditions. Among the most widely used estimators for heavy-tailed data is the Hill estimator, originally introduced by Hill1975. Its consistency in dependent settings was first investigated by Resnick_Starica_1995, who analysed its behaviour under various dependence structures. Further work by Resnick_Starica_1997 established the consistency of the tail empirical measure for autoregressive processes with heavy-tailed innovations. The asymptotic normality of the Hill estimator for i.i.d. data has been well documented, with significant contributions from Hall_1982Davis_Resnick_1984Haeusler_Teugels_1985, and Csorgo_Mason_1985. Extending these results to stationary sequences, HSING1991117 proved a central limit theorem for the Hill estimator under $\beta$-mixing conditions, and a modern treatment can be found in Kulik2020.
@@
Other work is related to ours. In the presence of covariates, the asymptotic properties of extreme value estimators have been investigated using regression-based approaches. In Daouia2010 tail index estimation is studied under $\alpha$-mixing conditions using a smoothed Pickands estimator, a technique later generalized by Daouia_2013; Daouia_Stupfler_Usseglio-Carleve_2023 to regression models within the general max-domain of attraction. Furthermore, tail index regression in a random covariate setting has been analysed using a maximum likelihood framework, as explored by Wang01092009. Finally, asymptotic normality of the conditional kernel Hill estimator has been explored through different techniques and for a special second-order regularly varying class in the preprint GoegebeurGuillou2024.
@@ 2.1 Notation and setting
Let $x\in\mathbb{R}$ with $g(x)>0$ be given. Our key assumption is that the conditional distribution of $Y_{0}$, given $X_{0}=x$, is regularly varying with tail index $1/\gamma(x)>0$, and that this holds in a neighbourhood around $x$, that is
@@
Let $K$ be a bounded density, denoted a kernel, supported on [1,1][-1,1] and denote by $\left\{h_{n},n\in\mathbb{N}\right\}$ the corresponding bandwidth (scale) parameter. Let $\left\{u^{x}_{n},n\in\mathbb{N}\right\}$ be a scaling sequence diverging to infinity. The conditional survival function $\overline{F}^{x}$ is estimated with the Nadaraya–Watson pathwise estimator as
$\overline{F}^{x}_{n}(y)=\sum_{j=1}^{n}\frac{K\left(\frac{x-X_{j}}{h_{n}}\right)}{\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)}1_{\left\{Y_{j}>y\right\}},\quad y>0.$
@@
$\widetilde{T}^{x}_{n}(s)=\frac{\overline{F}^{x}_{n}(su_{n}^{x})}{\overline{F}^{x}(u_{n}^{x})}=\frac{\frac{1}{nh_{n}\overline{F}^{x}(u_{n}^{x})}\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>su_{n}^{x}\right\}}}{\frac{1}{nh_{n}}\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)}=:\frac{\widetilde{V}^{x}_{n}(s)}{g_{n}(x)},$
for $s>0$. The target of $\widetilde{T}^{x}_{n}$ is the conditional tail function $s\mapsto T^{x}(s)=s^{-1/\gamma(x)}$. Indeed, this convergence, both in probability and weakly, will be shown under suitable conditions in this paper. However, the deterministic thresholds $u_{n}^{x}$ are not observable. Below we provide asymptotic results also for when these thresholds are estimated from data.
@@ 2.2 Conditional tail function and conditional Hill estimators
$\gamma(x)=\int_{1}^{\infty}s^{-1/\gamma(x)-1}\mathrm{d}s=\int_{1}^{\infty}\frac{T^{x}(s)}{s}\mathrm{d}s.$
@@
Let $u_{n}^{x}=F^{\leftarrow,x}\left(1-k_{n}/n\right)$ be an unobserved sequence of deterministic thresholds, diverging to infinity, where $F^{\leftarrow,x}$ denotes the generalised inverse of $F^{x}$. We may clearly estimate this sequence by its empirical counterpart $\left\{q_{n,k_{n}}^{x},n\in\mathbb{N}\right\}$ given by
$q_{n,k_{n}}^{x}=F^{\leftarrow,x}_{n}\left(1-k_{n}/n\right),\quad n\in\mathbb{N}.$
@@
$\widehat{T}^{x}_{n}(s)=\frac{\overline{F}^{x}_{n}(sq_{n,k_{n}}^{x})}{\overline{F}^{x}(u_{n}^{x})}=\widetilde{T}^{x}_{n}\left(s\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}\right),$
which no longer depends on the sequence $u_{n}^{x}$, which can be seen from the definition of $\widetilde{T}^{x}_{n}$ and from the fact that $\overline{F}^{x}(u_{n}^{x})=k_{n}/n$.
@@
$\widehat{\gamma}_{n}\left(x\right)=\int_{1}^{\infty}\frac{\widehat{T}_{n}^{x}(s)}{s}\mathrm{d}s=-\int_{1}^{\infty}\log\left(s\right)\widehat{T}_{n}^{x}(\mathrm{d}s)=\frac{n}{k_{n}}\frac{\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{j}}{q_{n,k_{n}}^{x}}\right)}{\sum_{j=1}^{n}K\left(\frac{x-X_{j}}{h_{n}}\right)}.$
@@
In case of $K$ being the uniform density on [1,1][-1,1], then $\widehat{\gamma}_{n}\left(x\right)$ reduces to the Hill estimator applied to the subsample of the data having $X_{i}\in[x-h_{n},x+h_{n}]$. Notice that the $k_{n}$ sequence no longer corresponds to the number of top order statistics in the subsample. Instead, let that number be defined as $\tilde{k}_{n}$. They are linked through the following relation:
@@
Figure 1: Top panel: covariate process $\left\{X_{j}\right\}$ given by the Crude Oil Volatility Index (CBOE OVX). The effective sample is calculated using a uniform kernel centred at x=0.05, 0.50, 0.95x=0.05,\,0.50,\,0.95 (highlighted in red, green and blue, respectively) and with $h_{n}=0.05$. Middle panel: target process $\left\{Y_{j}\right\}$ given by the absolute value of the negative log-returns of the West Texas Intermediate (WTI) grade crude oil spot prices with highlighted subsamples. Bottom panel: conditional Hill estimates $\widehat{\gamma}_{n}\left(x\right)$ in dotted blue, dashdotted green and dashed red, respectively; in solid grey we have the unconditional Hill estimator of the entire sample.
Before delving into the mathematical treatment of our estimators, we provide an example of a conditional Hill plot. In Figure 1 we have taken $\left\{Y_{j}\right\}$ to be the (absolute value of) the negative log-returns of the West Texas Intermediate (WTI) grade crude oil spot prices from 20072007-$2024$. As a covariate $\left\{X_{j}\right\}$, we take the Crude Oil Volatility Index (CBOE OVX), transformed to have uniform marginals. For large $x$, we clearly observe a different, heavier tail regime. In contrast, for medium and small $x$, the tail regime is similar and lighter. Below, it’s shown that the choice of kernel enters the asymptotic variance of the conditional Hill estimator through a certain integral. Transforming the covariate process to have uniform marginals is not necessary. It’s nonetheless done here, in the simulations and real data analysis below to avoid sparse regions of data to ensure good practical performance.
@@ 3 Main results
In this section we provide consistency and weak convergence results for the estimators of the conditional tail function and conditional tail index under suitable conditions. Most conditions are versions of their most general unconditional counterparts. Notably, only $\alpha$-mixing is required for consistency, while $\beta$-mixing is required for normality. Furthermore, second order conditions are formulated in a very general and abstract form. It is later seen that broad classes of regularly-varying tails satisfy them. Unless otherwise specified, all statements below hold for every $x\in U$.
@@ 3.1 Consistency
    The map $z\mapsto 1/\gamma(z)$ is $c_{\gamma}$-Lipschitz continuous,
@@
    $\lim_{n\to\infty}nh_{n}\overline{F}^{x}(u_{n}^{x})=\infty$,
@@
    It holds $g_{n}(x)\overset{\mathrm{Pr}}{\to}g(x)$, see Lemma 1213 or 14,
@@
    $\frac{\mathbb{E}\left[1_{\left\{Y_{0}>su_{n}^{x}\right\}}1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\mid X_{0}=x+h_{n}w_{1},X_{j}=x+h_{n}w_{2}\right]}{\overline{F}^{x}(u_{n}^{x})}\to M^{x}(s,t),$
@@
$\widetilde{T}^{x}_{n}(s)\overset{\mathrm{Pr}}{\to}T^{x}(s)=s^{-1/\gamma(x)},\ s>0,$ (2)
@@
If there exists $c_{x}\in(0,\infty)$ such that
$\frac{\overline{F}^{x}(u_{n}^{x})+h_{n}^{c_{x}\eta-2}}{nh_{n}^{c_{x}}(\overline{F}^{x}(u_{n}^{x}))^{2}}\to 0,$ (3)
@@
As a first application we get that consistency of the conditional tail empirical function implies that the intermediate empirical quantile $q_{n,k_{n}}^{x}$ is equivalent to the deterministic $u_{n}^{x}$. This is of practical importance, since the former is estimable from data. Consistency of the conditional Hill estimator then directly follows.
@@
If (2) holds, then $q_{n,k_{n}}^{x}/u_{n}^{x}\overset{\mathrm{Pr}}{\to}1.$
@@
Assume conditions 1-35 and (2). Then $\widehat{\gamma}_{n}\left(x\right)\overset{\mathrm{Pr}}{\to}\gamma(x)$.
@@ 3.2 Central limit theorems
Notice that the speed of convergence is determined by the effective sample size $nh_{n}\overline{F}^{x}(u_{n}^{x})=h_{n}k_{n}$, which is slower than in the unconditional case, which is simply $k_{n}$. This price to pay, however, results in a variance which agrees with the i.i.d. conditional case, in contrast to the more involved expression for unconditional time series provided as an infinite sum, e.g. in Kulik2020.
@@
$B_{n}^{x}(s_{0})=\sup_{t\geq s_{0}}\left|\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-1/\gamma(x)}g(x)\right|.$
@@
    For all $s,t>0$,
    $\lim_{m\to\infty}\limsup_{n\to\infty}\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)1_{\left\{Y_{0}>su_{n}^{x}\right\}}K\left(\frac{x-X_{j}}{h_{n}}\right)1_{\left\{Y_{j}>tu_{n}^{x}\right\}}\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}=0,$
@@
    For all $A>0$,
    $\lim_{m\to\infty}\limsup_{n\to\infty}\sum_{j=m}^{r_{n}}\frac{\mathbb{E}\left[K\left(\frac{x-X_{0}}{h_{n}}\right)K\left(\frac{x-X_{j}}{h_{n}}\right)\log_{+}\left(\frac{Y_{0}}{u_{n}^{x}A}\right)\log_{+}\left(\frac{Y_{j}}{u_{n}^{x}A}\right)\right]}{h_{n}\overline{F}^{x}(u_{n}^{x})}=0,$
@@
    $\lim_{n\to\infty}\sqrt{k_{n}h_{n}}\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-1/\gamma(x)}g(x)}{t}\mathrm{d}t=0.$
@@
$\Gamma(s,t)=\frac{\left(s\vee t\right)^{-1/\gamma(x)}}{g(x)}\int K^{2}(u)\mathrm{d}u,$
@@
$\sqrt{h_{n}k_{n}}\left\{\frac{q_{n,k_{n}}^{x}}{u_{n}^{x}}-1\right\}\overset{d}{\to}\gamma(x)\mathbb{T}^{x}(1),$
@@
$\gamma(x)\mathbb{T}^{x}(1)\overset{d}{=}\text{N}\left(0,\frac{\gamma(x)^{2}}{g(x)}\int K^{2}(u)\mathrm{d}u\right).$
The sequence $\{u_{n}^{x}\}$ is unobserved and we may estimate it with $\{q_{n,k_{n}}^{x}\}$. Let
@@
$\widehat{\mathbb{T}}^{x}_{n}\overset{w}{\Longrightarrow}\mathbb{T}^{x}-\mathbb{T}^{x}(1)\times T^{x},$
@@
$\sqrt{k_{n}h_{n}}\left(\widehat{\gamma}_{n}\left(x\right)-\gamma(x)\right)\overset{d}{\to}\text{N}\left(0,\frac{\gamma(x)^{2}}{g(x)}\int K^{2}(u)\mathrm{d}u\right).$
@@
provided conditions involving the logarithmic function are replaced with suitably modified versions involving $\phi$. For instance, the choice $\phi(s)=\log(s)^{2}$, yields an estimator of $2\gamma(x)^{2}$. In case the functional $z\mapsto\int\phi\mathrm{d}y$, or any other functional is defined on $\mathbb{D}\left([A,B]\right)$ and is sup-norm continuous or Hadamard differentiable, consistency and a central limit theorem for the associated estimator follow directly from Theorem 4, together with the continuous mapping theorem or functional delta-method, respectively.
Remark.
The bias of the conditional Hill estimator vanishes due to condition 7. If this condition is relaxed to
$\lim_{n\to\infty}\sqrt{k_{n}h_{n}}\int_{1}^{\infty}\frac{\mathbb{E}\left[\widetilde{V}^{x}_{n}(t)\right]-t^{-1/\gamma(x)}g(x)}{t}\mathrm{d}t\ =\ B\in\mathbb{R},$
the asymptotic variance in Theorem 5 remains similar, while the bias becomes $B/g(x)$.